If m<1=9x+6, m<2=2(5x-3),and m<3=5y+14. Find X and Y. I can find X but not sure about Y as it doesn't look right.
Answer by ikleyn(52790)

.
If m<1=9x+6, m<2=2(5x-3),and m<3=5y+14. Find X and Y.
I can find X but not sure about Y as it doesn't look right.
~~~~~~~~~~~~~~


Who are these personages X and Y in your post,

and how are they connected with the other personages in your post ?


Why do you use capital and lowercase x an y --- are they different ?

Are these angles   < 1,   < 2   and   < 3   the angles of a triangle ?   Or some other figure ?



Give the problem condition/description in full and in precise terms.